Suppose you wanted to use a non-reflecting layer for radar waves to make an aircraft invisible. What would the thickness of the layer be to avoid reflecting 2 cm radar waves. (You can neglect changes of wavelength in the layer for this problem.) Would there be any problems as the aircraft turn? Can you imagine any other ways to use the wave nature of radar to try to make "stealth" aircraft? How could an on-board transmitter help?
Answer #1

For destructive interference

2t=\left ( m+\frac{1}{2} \right )\frac{\lambda}{n}

Where:

t: tichkness

lambda :2cm

n: depending of the material--> we can asume 1.6

m=0; the most thick

clearing t and substituting;

t=\frac{\lambda}{4n}=\frac{2cm}{4*1.6}=0.3125cm

however this depend of the index of refraction of the material (n).

Would there be any problems as the aircraft turn?

no. the radar can receive only the waves normally inciident. and this waves reflected will be destroyed. so, when the aircraft turn the incident waves always be destroyed.

Can you imagine any other ways to use the wave nature of radar to try to make "stealth" aircraft?

there are materials that can absorb the energy of the waves and transform in heat, reducing in this way the reflected wave. this materials are knowed like radiation absorbent materials. if the aircraft is covered by this materials this could help to be invisible to some frecuencies of radar.

How could an on-board transmitter help?

with precision a transmitter could send a signal of the same frecuency that the signal of the radar and with a phase destructive (50%) and so, be indetectable



answered by: ANURANJAN SARSAM